# Open, Monetizable, and Loyal AI

### v1

@@ 4.2.4 Incentive Layer
Usage Tracking and Payment
@@
4.3 Blockchain for Transparency and Trust
A programmable blockchain technology [120] offers a transparent, immutable, and trust-minimized infrastructure that allows natural implementations for the designs in the Sentient Protocol of the previous section. By integrating such blockchain components across specific layers of the protocol, the requirements for openness, trust-free, and fairness are readily met efficiently and securelhy. A summary of which modules of the Sentient layered design are implemented on blockchain is presented in Figure 4.6.
@@ 4.3 Blockchain for Transparency and Trust
Figure 4.6: The dashed line shows where the blockchain stands in each of the layers of the Sentient protocol.
Blockchain in the Storage Layer
@@
The evaluator module needs a mechanism to prove the result of the evaluation “on-chain”, so that anyone can verify that the model is evaluated correctly, and hence the ownership allocation can be trusted. Standard benchmarks together with various “proofs of inference” are a natural solution [122].
Blockchain in the Access Layer
@@
Blockchain in the Incentive Layer
@@
4.4 A Sentient Protocol Implementation of OML 1.0
@@
In this section, we combine the fingerprinting-based OML mechanism of Chapter 3 with the incentives and architectural framework of the Sentient protocol; tying them together results in a Sentient protocol implementation of OML 1.0.
@@
Under this paradigm, upon completing a new model M, model builders first submit this model to the Sentient protocol distribution layer which securely stores this raw model and tracks which builders own what stake in this model - this will correspond to how much payment they are due when the model is in use. When a model user (who hosts the model for services to external end users) requests to access the model, converter nodes within the distribution layer convert the model into a unique OMLized model, $M$.oml for that user. This conversion process implements OML 1.0 by injecting several secret (key, response) pairs that fingerprint the model unique to the model user. The model is trained on these fingerprint pairs such that for any input of a key, the OMLized model will output a corresponding secret response (see Figure 4.7). These key response pairs can be highly diverse in nature, random or structured, insertable in the thousands, and together very resilient against fine-tuning or model modification. Chapter 3 discusses our research on this technique and its limits in depth.
@@
Figure 4.7: An $M$.oml model is fine-tuned on numerous secret fingerprint (key, response) pairs, that can later be used to track the provenance of the model.
@@
From here, this model is distributed to model users. Note that the term user does not exclusively refer to final end users but also include model hosts, such as a company hosting a chatbot or tool powered by a Sentient model, which is then served via API or a web interface to final end users. In order for the model user to access the OMLized model as outlined above, the model user must provide adequate collateral in the form of a deposit made to the access layer of the protocol, which also tracks which models are distributed to which model users. This collateral grants a crypto-economic barrier to the model user using or distributing the model without complying with license terms – payment – to the model builders (see Figure 4.8).
@@
Figure 4.8: Model builders submit raw models to the Sentient Protocol. Model users then deposit collateral with the protocol to access a unique fingerprinted M.oml model.
@@
Model users understand that they are required to request permission from the access layer and make any corresponding payment, when fulfilling any external requests by an end user to run inference on the OMLized model. In the case that a prover can prove that a model user has not done so, the model user loses some or all of the collateral they have deposited with the protocol (see Figure 4.9 presenting correct usage of the protocol).
@@
Figure 4.9: When an end user submits a query to a model user (via a web application or API for example), model users pass a permission request and payment for this usage of the OML model to the Sentient protocol.
@@
Figure 4.10: Provers verify that a model user is complying with the protocol by posing as a benign end user and querying the model user (via a web application or API for example) with one of the secret fingerprint keys they hold. The prover can submit the key and the user’s response to the protocol, which can then in turn determine if the model user appropriately requested permission for the request.
If the model user appropriately requested permission, no further action is required as the protocol is functioning as it is intended to. If the model user is demonstrated to have used the model without requesting permission, however, the model user can be penalized via the collateral that they were required to post when they received the model in the first place. Note that if the model user shares the model with a third party who then illegally uses the model, the original model user is penalized for any misuse by the third party (see Figure 4.10).
As noted in Section 3.4, the above paradigm can be extended beyond monetization to enforce some degree of loyalty of AI models as well. Now the provers also check if model users are modifying the model to mitigate safety measures implemented in training. OMLization would now involve modifying the fine-tuning dataset to add some prompts with safe responses. Provers could use the more robust (key, response) pairs to identify the provenance of the model while testing known safe key response pairs to see if model users have tampered with model safeguards.
4.5 Concluding Remarks
In the era of AI entrepreneurship, individual contributors of AI models (and other artifacts) will be owners of the artifacts they help build. There are two technical challenges that need to be addressed in enabling AI ownership:
@@ 4.5 Concluding Remarks
A complete implementation of evaluator and ownership modules requires us to choose different methods for evaluating contributions to AI models and datasets, executing these evaluations under different trust models, and using different governance contracts for assigning ownership based on evaluated contribution. These details are left to a future version of this paper. Details on how the queries will be shared with the watchers, and the nuances of proving as a bounty hunter watcher, will be discussed in a future version, too.
Contributors
The contributors are listed in alphabetical order by last name. The authors can be contacted via the email address firstname@@@@sentient.foundation.
  • Zerui Cheng. Zerui was a research intern at Sentient; he is pursuing his PhD at Princeton University.
  • Edoardo Contente. Edoardo is a researcher at Sentient; he recently graduated with A.B. and M.Eng degrees from Princeton University.
  • Ben Finch. Ben is a researcher at Sentient; he recently graduated with B.S.E. and M.Eng degrees from Princeton University.
  • Oleg Golev. Oleg is a researcher at Sentient; he recently graduated with B.S.E. and M.Eng degrees from Princeton University.
  • Jonathan Hayase. Jon is a research intern at Sentient while pursuing his PhD at the University of Washington.
  • Andrew Miller. Andrew is a collaborating researcher at Sentient; he is a professor at the University of Illinois at Urbana-Champaign.
  • Niusha Moshrefi. Niusha was a research intern at Sentient; she is pursuing her PhD at Princeton University.
  • Anshul Nasery. Anshul is a research intern at Sentient while pursuing his PhD at the University of Washington.
  • Sandeep Nailwal. Sandeep is a core contributor at Sentient; he is a co-founder of Polygon.
  • Sewoong Oh. Sewoong is a researcher at Sentient; he is a professor at the University of Washington.
  • Himanshu Tyagi. Himanshu is a core contributor at Sentient; he is a professor at the Indian Insitute of Science.
  • Pramod Viswanath. Pramod is a core contributor at Sentient; he is a professor at Princeton University.
Corresponding author is Pramod Viswanath. Email: pramod@@@@sentient.foundation.

### v3

@@ C.2.4 Incentive Layer
Usage Tracking and Payment.
@@
C.3 Blockchain for Transparency and Trust
A programmable blockchain technology [108] offers a transparent, immutable, and trust-minimized infrastructure that allows natural implementations for the designs in the Sentient Protocol of the previous section. By integrating such blockchain components across specific layers of the protocol, the requirements for openness, trust-free, and fairness are readily met efficiently and securelhy. A summary of which modules of the Sentient layered design are implemented on blockchain is presented in Figure 18.
@@ C.3 Blockchain for Transparency and Trust
Figure 18: The dashed line shows where the blockchain stands in each of the layers of the Sentient protocol.
Blockchain in the Storage Layer.
@@
The evaluator module needs a mechanism to prove the result of the evaluation “on-chain", so that anyone can verify that the model is evaluated correctly, and hence the ownership allocation can be trusted. Standard benchmarks together with various “proofs of inference" are a natural solution [14].
Blockchain in the Access Layer.
@@
Blockchain in the Incentive Layer.
@@
C.4 A Sentient Protocol Implementation of OML 1.0
@@
In this section, we combine the fingerprinting-based OML mechanism of Appendix B with the incentives and architectural framework of the Sentient protocol; tying them together results in a Sentient protocol implementation of OML 1.0.
@@
Under this paradigm, upon completing a new model M, model builders first submit this model to the Sentient protocol distribution layer which securely stores this raw model and tracks which builders own what stake in this model - this will correspond to how much payment they are due when the model is in use. When a model user (who hosts the model for services to external end users) requests to access the model, converter nodes within the distribution layer convert the model into a unique OMLized model, $M$.oml for that user. This conversion process implements OML 1.0 by injecting several secret (key, response) pairs that fingerprint the model unique to the model user. The model is trained on these fingerprint pairs such that for any input of a key, the OMLized model will output a corresponding secret response (see Figure 19). These key response pairs can be highly diverse in nature, random or structured, insertable in the thousands, and together very resilient against fine-tuning or model modification. Appendix B discusses our research on this technique and its limits in depth.
@@
Figure 19: An $M$.oml model is fine-tuned on numerous secret fingerprint (key, response) pairs, that can later be used to track the provenance of the model.
@@
From here, this model is distributed to model users. Note that the term user does not exclusively refer to final end users but also include model hosts, such as a company hosting a chatbot or tool powered by a Sentient model, which is then served via API or a web interface to final end users. In order for the model user to access the OMLized model as outlined above, the model user must provide adequate collateral in the form of a deposit made to the access layer of the protocol, which also tracks which models are distributed to which model users. This collateral grants a crypto-economic barrier to the model user using or distributing the model without complying with license terms – payment – to the model builders (see Figure 20).
@@
Figure 20: Model builders submit raw models to the Sentient Protocol. Model users then deposit collateral with the protocol to access a unique fingerprinted M.oml model.
@@
Model users understand that they are required to request permission from the access layer and make any corresponding payment, when fulfilling any external requests by an end user to run inference on the OMLized model. In the case that a prover can prove that a model user has not done so, the model user loses some or all of the collateral they have deposited with the protocol (see Figure 21 presenting correct usage of the protocol).
@@
Figure 21: When an end user submits a query to a model user (via a web application or API for example), model users pass a permission request and payment for this usage of the OML model to the Sentient protocol.
@@
Figure 22: Provers verify that a model user is complying with the protocol by posing as a benign end user and querying the model user (via a web application or API for example) with one of the secret fingerprint keys they hold. The prover can submit the key and the user’s response to the protocol, which can then in turn determine if the model user appropriately requested permission for the request.
If the model user appropriately requested permission, no further action is required as the protocol is functioning as it is intended to. If the model user is demonstrated to have used the model without requesting permission, however, the model user can be penalized via the collateral that they were required to post when they received the model in the first place. Note that if the model user shares the model with a third party who then illegally uses the model, the original model user is penalized for any misuse by the third party (see Figure 22).
As noted in Section B.4, the above paradigm can be extended beyond monetization to enforce some degree of loyalty of AI models as well. Now the provers also check if model users are modifying the model to mitigate safety measures implemented in training. OMLization would now involve modifying the fine-tuning dataset to add some prompts with safe responses. Provers could use the more robust (key, response) pairs to identify the provenance of the model while testing known safe key response pairs to see if model users have tampered with model safeguards.
C.5 Concluding Remarks
In the era of community-built AI, individual contributors of AI models (and other artifacts) will be owners of the artifacts they help build. There are two technical challenges that need to be addressed in enabling AI ownership:
@@ C.5 Concluding Remarks
A complete implementation of evaluator and ownership modules requires us to choose different methods for evaluating contributions to AI models and datasets, executing these evaluations under different trust models, and using different governance contracts for assigning ownership based on evaluated contribution, and finalize details on how the queries will be shared with the watchers, and the nuances of proving as a bounty hunter watcher.